how do you convert a point slope for example y+3=2/5(x+10) into standard form
Answer by stanbon(75887)

how do you convert a point slope for example y+3=2/5(x+10) into standard form
===============
y+3=2/5(x+10)
------
y + 3 = (2/5)(x+10)
----
Multiply thru by 5 to get:
5y + 15 = 2x+20
----
Standard Form: ax + by = c
2x - 5y = -5
---------------------
Cheers,
Stan H.